# Evolving Symbolic 3D Visual Grounder with Weakly Supervised Reflection

### v2

@@ 3.3.1 Class Schema
For each relation, we define a corresponding Python class as its encoder. The class is initialized with the scene’s point cloud data and object segmentation and provides two key methods: _init_param, which computes the necessary parameters for feature derivation. For instance, in the “near” class, it calculates distances between each pair of objects; forward: which performs numerical operations such as inversion or exponentiation and returns the relation feature.
@@ 3.3.2 In Context Example
Adding in-context example (ICE) into prompt can significantly improve the accuracy of responses from LLM (Brown et al., 2020). To reduce human effort and provide suitable ICE for different relation encoders’ generation, we utilize a semantic based retrieval strategy. For example, relation encoders for “near" and “far" may both compute pairwise distances but differ only in the numerical processing, so the well optimized relation encoder for “near" can serve as ICE for generation of “far".
We construct a directed acyclic graph $G$ (see Figure 6). Each node in $G$ represents a specific relation, and an edge from node A to node B indicates that the encoder for relation A is used as an ICE when generating for relation B. With this structure, the entire code generation process follows a topological ordering over $G$ automatically.
@@ 3.3.3 Unit Test
LLMs may not always generate a perfect code within one attempt (Olausson et al., 2023). Wu et al. (2024) proposes that more sampling increases the likelihood of correct answers. So we design unit test suites to select the plausible relation encoders from multiple LLM responses.
Take binary relation “above" for example, we collect a few triplets in the format of target object, same class distractor, anchor object from training set, with each triplet serving as a unit test case. The scale for each relation is small (less than 50 for most of relations). In the relation feature from encoder, if the distractor-anchor pair (probability of distractor “above" anchor) is larger than the target-anchor pair, the test fails and an error message like [target bbox] is above [anchor bbox] So metric value of [target bbox] "above" [anchor bbox] should be larger than than the metric value of [distractor bbox] "above" [anchor bbox]. is synthesized. 3 shows an example of the error message.
@@ 3.3.4 Code Generation and Optimization
As illusrated in Figure 2, for any relation, we begin by prompting the LLM with a high-level task description, the relation name, and an in-context example retrieved from codes that are already optimized as §$\S$ 3.3.2. Then we sample $N_{\text{sample}}$ candidate code snippets from LLM, where $N_{\text{sample}}$ is a configurable hyperparameter.
Next, each generated code is tested using the unit tests defined in §$\S$ 3.3.2. We select the $top_{k}$ codes that pass the most test cases and subject them to an optimization phase. During optimization phase, the LLM is given the initial prompt, the code to be optimized, and error message produced by the test suite. It then revises the code according to these errors.
This procedure of testing and refinement is repeated for up to $N_{\text{iter}}$ iterations. Ultimately, we adopt the code that achieves the highest pass rate across all test cases. In the event of a tie, we select the code that underwent more optimization steps. The optimization and selection algorithm is shown in Algorithm 1.
@@
Hyperparameters : search iteration $N$, sample number $M$, evolving example number $K$
@@ 3.4 VLM Decision
The visual information like description on color or shape in utterances can be essential for accurate grounding, particularly for natural datasets like Nr3D (Achlioptas et al., 2020) and ScanRefer (Chen et al., 2020). When two candidate objects share similar class and spatial position, visual perception is required to distinguish between them.
Following Xu et al. , we incorporate GPT-4o to identify the target object from a set of candidates by utilizing 2D images from ScanNet (Dai et al., 2017) as additional context. Specifically, we select the top five objects based on output from executor and retain those whose logits exceed a chosen threshold as final candidates. Next, we retrieve the corresponding images from ScanNet (Dai et al., 2017) by matching each candidate’s point cloud to camera parameters, thereby finding images that include the candidate objects. Out of these, we select eight images that contain the most prominent views of each candidate (i.e., the largest bounding boxes in the image). We then annotate each image with object IDs and stitch them together in a $4\times 2$ grid. Finally, we prompt GPT-4o to identify the target object from stitched images. By integrating these visual cues, the VLM decision step effectively disambiguates candidates that appear similar in terms of class and spatial attributes, yielding more accurate results.
@@ 4.1 Experimental Settings
We conduct experiments on Nr3D subset from ReferIt3D (Achlioptas et al., 2020) dataset and ScanRefer (Chen et al., 2020). ReferIt3D has 2 subsets: Nr3D and Sr3D. Nr3D subset utterances contain human-annotated utterances and Sr3D contains synthesized ones. Based on the number of same-class distractors, the dataset can be categorized into “easy" and “hard" subsets depending on the number of same class distractors. The easy subset has a single distractor and the hard subset has multiple. The dataset can also be split into “view dependent" and “view independent" subsets according to if there are some keywords in the utterance. Ground truth object bounding boxes are given in ReferIt3D default test setting. So the metric is an exact match between the predicted bounding box and the target bounding box. In ScanRefer, no GT object mask is provided, the evaluation metric is the intersection over union (IoU) value between predicted bounding box and GT bounding box. There are two different thresholds: 0.25 and 0.5, so the metric names are Acc@0.25, Acc@0.5 respectively.
@@
For code evolving, we set $N_{sample}$ and $N_{iter}$ to 5, $topk$ to 3. We mainly use gpt-4o-2024-08-06 model with a temperature of 1.0 and top_p of 0.95. For a fair comparison, we use the same object classification result and evaluation code as  Yuan et al. (2024b) for evaluation of ReferIt3D. For evaluation of ScanRefer, we use the object detection and classification results of (Jain et al., 2024). (internal experiments show that EaSe has similar performance when using same object proposals as Yuan et al. (2024b).) For VLM decision, we use the same temperature and top_p as Xu et al. , the threshold is 0.9 for Nr3D, and 0.1 for ScanRefer.
@@
We compare EaSe against both supervised and training-free methods, evaluating their performance and inference cost. The supervised baselines include BUTD-DETR (Jain et al., 2022), Vil3DRel (Chen et al., 2022), 3D-VisTA (Ziyu et al., 2023), and CoT3DRef (Bakr et al., 2023), while the training-free approaches include ZSVG3D (Yuan et al., 2024b), VLM-Grounder (Xu et al., ), and Transcrib3D (Fang et al., 2024).
For the Nr3D dataset, Transcrib3D (Fang et al., 2024) uses ground-truth object labels, which are substantially more accurate than the predicted labels we employ. Additionally, the neuro-symbolic method NS3D (Hsu et al., 2023) operates under a more limited evaluation protocol; therefore, we compare to them within their specific settings.
@@ 4.2 Quantitative Results
Table 4 presents the results on Nr3D. Compared to other training-free baselines, EaSe achieves better overall accuracy than both ZSVG3D (Yuan et al., 2024b) and VLM-Grounder (Xu et al., ). When combined with VLM, EaSe further narrows the gap in overall performance relative to the supervised BUTD-DETR (Jain et al., 2022), especially on the view-dependent (VD) subset. However, it still lags behind other more recent supervised methods, which are trained on large scaled specific 3D datasets. We evaluate EaSe in the settings of Fang et al. (2024) and Hsu et al. (2023) respectively. In Fang et al. (2024)’s setting, the GT object labels are utilized for more accurate category-level object recognition. EaSe underperforms Fang et al. (2024) by 2.4% (67.8% vs. 70.2%). The high accuracy of Fang et al. (2024) comes from LLM based reasoning during inference time, which greatly hurts its inference efficiency. In the setting of  Hsu et al. (2023), EaSe outperforms it by 7.4% without any training.
@@
Table 3 shows the results on ScanRefer. VLM-Grounder (Xu et al., ) is evaluated on a randomly selected subset containing 250 samples EaSe outperforms ZSVG3D (Yuan et al., 2024b) by $12.8\%$ (49.2% vs. 36.4%) when leveraging a VLM. Moreover, the performance gap between EaSe (49.2%) and Xu et al. (51.6%), Fang et al. (2024) (51.3%) is relatively small, demonstrating that EaSe balances efficiency with competitive accuracy.
@@
Table 2 compares the inference time and token consumption of various training-free methods on Nr3D. We randomly sampled a subset of the Nr3D test set to calculate the average inference time and token usage. Agent-based methods (Xu et al., ; Fang et al., 2024) rely on multiple calls to the LLM during a single inference, with the context continuously growing. As a result, these methods consume significantly more time and tokens (27.0s and 50.3s, respectively) compared to EaSe (2.1s).
In contrast, methods like EaSe (without VLM) and ZSVG3D (Yuan et al., 2024b) only require a single call to the LLM for semantic parsing with a short prompt, leading to lower time and token consumption during inference, and EaSe significantly outperforms ZSVG3D in accuracy by a large margin. When combined with a VLM, EaSe still achieves more than three times faster inference and significantly reduced token consumption compared to agent-based methods (Xu et al., ; Fang et al., 2024).
In conclusion, the above quantitative results highlight that EaSe effectively balances performance and efficiency, achieving leading accuracy alongside exceptional computational cost savings among training-free methods.
Table 2: Grounding time and token cost on Nr3D for training-free method that utilized LLM. EaSe has significant advantage, especially when compared to agent based methods (VLM-Grounder and Transcrib3D). $\dagger$: evaluated on a subset having 250 samples.
@@
Table 3: Performance on ScanRefer. EaSe has close performance with VLM-Grounder and BUTD-DETR. $\dagger$: evaluated on a subset having 250 samples.
@@
Table 4: Performance on Nr3D. VD and VID stand for view-dependent and view-independent, respectively. We only report time and token consuming for training-free methods which involve LLM or VLM. EaSe outperforms other training-free baselines in overall performance, and has significant advantages in terms of time and token consumptions. For comparison with Transcrib3D, we use GT object labels. For comparison with NS3D, we evaluate EaSe on the same Nr3D subset as NS3D. $\dagger$: VLM-Grounder is evaluated on a subset having 250 samples. * : we re-run ZSVG3D using GPT-4o.
@@ 4.3 Qualitative Results
Figure 3: Explanation and visualization of the grounding steps. Anchors (desk for the upper case and door for the lower) are marked by red circle. We only visualize a part of objects that match the below conditions well in green, objects of brighter color have higher scores of meeting the condition.
@@
In Figure 3, we present two step-by-step grounding process of EaSe , illustrating how the final grounding results are constructed through combinition of multiple conditions within the referring utterance. The corresponding utterance for the second row is, “When facing the door, it’s the shelf above the desk on the right." The execution of the symbolic expression involves four steps, progressing from left to right. Objects with higher scores are highlighted in green, where lighter green indicates closer match between an object the condition below. First, we can find objects “right of the door" by the category feature of “door" and relation feature “right", which are in green color. Then we can find “desk right of the door" from them using the category feature of “desk". After that, objects which satisfied “above desk and on door’s right" can be found by joint condition of “above desk" and “right of the door". Finally the target “shelf" is grounded from them. More visualization results can be found in the Figure 8.
4.4 Ablation Studies
Figure 4: The learning curves of different variants. The x-axis is the generation number of the program. The y-axis is the accuracy of the program on the test set. It is worth noting that the curves of our full model Refer to caption coincide with the curves of Variant 3 Refer to caption in the second row (“Left”, “Above”, and “Corner”) since Variant 3 can be regarded as the first stage of generating all programs.
@@ 4.4 Ablation Studies
We conduct ablation studies to investigate the impact of various components during the code generation process, evaluating three different variants. We chose to analyze six relations that are relatively complex and have multiple iterations during the generation and optimization process. (For relatively simple relations like “small", the generated code passes all unit tests in the first generation so no optimization is used.) In all three variants, no in-context example (§$\S$3.3.2) is provided. The three different variants are: 1) Generation only: In this variant, we generate the code by directly prompting the LLM and sample multiple codes. The code with the highest pass rate on unit tests is selected. 2) Self-optimization only: In this variant, we use the same prompt as variant 1) and replace the feedback message (§$\S$3.3.4) containing specific failure cases with a general optimization instruction. 3) No ICE: This variant has feedback message §$\S$3.3.4 and no in-context example is used for generation. For the relations that no in-context examples available(“left”, “above”, and “corner”), variant 3 is identical to the first stage of full version of EaSe , so we only plot variants 1 and 2 on the corresponding plots. For relations “right", “below" and “between", we conduct experiment on all 3 variants. To control for the impact of the initial generation, we use the same results of iteration 0 across variant 2 and variant 3.
Figure 4 illustrates the results of ablation study, different variants are represented by lines of different colors. The horizontal axis represents the number of iterations, after an iteration ends, we choose a code based on their unit test pass rates and then evaluate it on the test set. The vertical axis shows the number of correctly solved test examples associated with the relation. We normalize them by dividing them by the maximum value. We can see that random sampling (Variant 1) shows comparable performance to the full model only for the “corner" and “between" relations. In Variant 2, which uses simple self-reflection, there is noticeable performance improvement across most relations except "corner." This variant achieves performance on "above" and "corner" close to that of the full model, as these relations are easier to generate. However, it struggles with the more complex relations. Variant 3, which incorporates feedback messages, improves upon the performance of Variant 2. For all relations except “right", it achieves results similar to the full model. However, for “right," “between," and “below" — where in-context examples are used to assist generation — Variant 3 shows a significant performance gap in the early iterations compared to the full model, highlighting the impact of in-context examples.
@@
In this work, we propose EaSe , a training free method that uses Python code as relation encoder for 3DVG and an automatic pipeline for code generation and optimization. EaSe does not rely on large-scale data and has promising advantages in performance, and inference cost compared to other similar methods. We directly use knowledge LLM for relation encoder instead of human annotation or supervised learning.
However, some limitations still remain. In referring utterances, there are diverse referring utterances containing words other than objects and relations. But our method mainly focuses on objects and spatial relations so it cannot handle the non-relation words in some complex cases. Besides, we treat ignore the objects’ orientations, and global information of the scene.

### v3

@@ 3.3.1 Class Schema
EaSe uses Python classes as spatial relation encoders. A class is initialized with the scene’s point cloud data and object segmentation and provides two key methods: _init_param, which computes the necessary parameters for feature derivation. For instance, in the “near” class, it calculates distances between each pair of objects; forward, which performs numerical operations such as inversion or exponentiation and returns the relation feature.
@@ 3.3.2 In Context Example
Adding in-context examples into the prompt can significantly improve the accuracy of responses from LLMs (Brown et al., 2020). To reduce human effort and provide suitable ICE for different relation encoders’ generation, we introduce a semantic-based retrieval strategy. For example, relation encoders for “near” and “far” may both compute pairwise distances but differ only in the numerical processing, so the well-optimized relation encoder for “near” can serve as an in-context example for the generation of “far”. We show the retrieval details in Figure 6.
@@ 3.3.3 Unit Test
Since LLMs may not always generate correct code in a single attempt (Olausson et al., 2023), inspired by Wu et al. (2024)’s finding that increased sampling enhances success probability, we design unit test suites to select the plausible relation encoders from multiple LLM responses.
Take the binary relation “above” as an example. We collect a few triplets in the format of target object, same class distractor, anchor object from the training set, with each triplet serving as a unit test case. The scale for each relation is small (less than 50 for most relations). In the generated relation feature for “above” $f$, if $f^{(distractor,anchor)}$ is larger than $f^{(target,anchor)}$, the test is deemed to have failed, and an error message like [target bbox] is above [anchor bbox] So metric value of [target bbox] "above" [anchor bbox] should be larger than the metric value of [distractor bbox] "above" [anchor bbox]. is synthesized. 3 provides an example of such an error message.
@@ 3.3.4 Code Generation and Optimization
As illustrated in Figure 2, for any relation, we begin by prompting the LLM with a high-level task description, the relation name, and an in-context example retrieved from codes that are already optimized as §$\S$ 3.3.2. Then we sample $N_{\text{sample}}$ candidate encoders from LLM, where $N_{\text{sample}}$ is a configurable hyperparameter. Next, each generated code is tested using the unit tests defined in § 3.3.3. We select the $top_{k}$ codes that pass the most test cases and subject them to an optimization phase. During the optimization phase, the LLM receives the initial prompt, the code to be optimized, and the error message produced by the test suite. It then revises the code according to these errors. This test and optimization procedure is repeated for up to $N_{\text{iter}}$ iterations. Ultimately, we adopt the code that achieves the highest pass rate across all test cases. In the event of a tie, we select the code that underwent more optimization steps. The optimization and selection algorithm is shown in Algorithm 1.
@@
Hyperparameters : search iteration $N$, sample number $M$, optimizing example number $top_{k}$
@@ 3.4 VLM Decision
The visual information, like descriptions on color or shape in utterances, can be essential for accurate grounding, particularly for natural datasets like Nr3D (Achlioptas et al., 2020) and ScanRefer (Chen et al., 2020). When two candidate objects share a similar class and spatial position, visual perception is required to distinguish between them.
Following Xu et al. , we incorporate GPT-4o to identify the target object from a set of candidates by utilizing 2D images from ScanNet (Dai et al., 2017) as additional context.
Specifically, we select the top five objects based on the scores from the relation encoders and retain those whose logits exceed a chosen threshold as final candidates. Next, we retrieve the corresponding images from ScanNet (Dai et al., 2017) by matching each candidate’s point cloud to camera parameters, thereby finding images that include the candidate objects. Out of these, we select eight images that contain the largest projected bounding box area of each candidate. We then annotate each image with object IDs and stitch them together in a $4\times 2$ grid. Finally, we prompt GPT-4o to identify the target object from stitched images. By integrating these visual cues, the VLM decision step effectively disambiguates candidates that appear similar in terms of class and spatial attributes, yielding more accurate results.
@@ 4.1 Experimental Settings
We conduct experiments on the Nr3D subset of ReferIt3D (Achlioptas et al., 2020) dataset and ScanRefer (Chen et al., 2020). ReferIt3D has 2 subsets: Nr3D and Sr3D. The Nr3D subset utterances contain human-annotated utterances and the Sr3D contains synthesized ones. Based on the number of same-class distractors, the dataset can be categorized into “easy” and “hard” subsets. The easy subset has a single distractor, and the hard subset has multiple distractors. The dataset can also be split into “view dependent” and “view independent” subsets according to the referring utterance. Ground truth object bounding boxes are given in the ReferIt3D default evaluation setting. Therefore, the metric is an exact match between the predicted bounding box and the target bounding box. In ScanRefer, no GT object mask is provided; the evaluation metric is the intersection over union (IoU) value between the predicted bounding box and the GT bounding box. We use the Acc@0.25 metric here.
@@
For code optimization (§ 3.3.4), we set $N_{sample}$ and $N_{iter}$ to 5, $top_{k}$ to 3. We mainly use gpt-4o-2024-08-06 model with a temperature of 1.0 and top_p of 0.95. For a fair comparison, we use the object classification results from  Yuan et al. (2024b) for the evaluation of ReferIt3D. For evaluation of ScanRefer, we use the object detection and classification results of (Jain et al., 2024). (Experiments show that EaSe has similar accuracy when using object detection and classification results from Yuan et al. (2024b).) For VLM decision-making, we use the same temperature and top_p values as in Xu et al. . The threshold (§ 3.4 is set to 0.9 for Nr3D and 0.1 for ScanRefer.
@@
We compare EaSe against both supervised and training-free methods, evaluating accuracy, grounding time, and token cost. The supervised baselines include BUTD-DETR (Jain et al., 2022), Vil3DRel (Chen et al., 2022), 3D-VisTA (Ziyu et al., 2023), and CoT3DRef (Bakr et al., 2023), while the training-free approaches include ZSVG3D (Yuan et al., 2024b), VLM-Grounder (Xu et al., ), and Transcrib3D (Fang et al., 2024).
On the Nr3D dataset, Transcrib3D (Fang et al., 2024) uses ground-truth object labels, providing an advantage over methods which rely on predicted labels. Additionally, the neuro-symbolic method NS3D (Hsu et al., 2023) operates under a more limited evaluation protocol; therefore, we compare EaSe with them under their specific settings.
@@ 4.2 Quantitative Results
Table 4 presents the results on Nr3D. Compared to other training-free baselines, EaSe (without VLM) achieves higher overall accuracy than both ZSVG3D (Yuan et al., 2024b) and VLM-Grounder (Xu et al., ). When integrated with VLM, EaSe further narrows the gap in overall performance relative to the supervised BUTD-DETR (Jain et al., 2022), especially on the view-dependent (VD) subset. However, it still lags behind other more recent supervised methods, which are trained on large-scale task-specific 3D datasets. We further evaluate EaSe under the experimental settings of Fang et al. (2024) and Hsu et al. (2023) respectively. In Fang et al. (2024)’s setting, the ground truth (GT) object labels are utilized for more accurate category-level object recognition. EaSe slightly underperforms Fang et al. (2024) by 2.4% (67.8% vs. 70.2%), but EaSe has significant advantages in grounding time and token cost. Under the setting of  Hsu et al. (2023), EaSe outperforms it by 7.4% without requiring any training.
@@
Table 3 shows the results on ScanRefer. EaSe outperforms ZSVG3D (Yuan et al., 2024b) by $12.8\%$ (49.2% vs. 36.4%). Moreover, the performance gap between EaSe (49.2%) and Xu et al. (51.6%), Fang et al. (2024) (51.3%) is relatively small, demonstrating that EaSe balances cost with competitive accuracy.
@@
Table 2 compares the average grounding time and token costs of various training-free methods on a randomly sampled subset of the Nr3D dataset. Agent-based methods (Xu et al., ; Fang et al., 2024) exhibit significantly higher time and token consumption (27.0s and 50.3k tokens, respectively) compared to EaSe (without VLM) (2.1s and 3.2k tokens). EaSe (without VLM) and ZSVG3D (Yuan et al., 2024b) have much lower costs and EaSe demonstrates a significant improvement in accuracy over ZSVG3D (Yuan et al., 2024b). Even when integrated with a VLM, EaSe maintains a more than threefold reduction in grounding time and token consumption compared to agent-based methods (Xu et al., ; Fang et al., 2024).
These quantitative results underscore the ability of EaSe to effectively balance performance and efficiency, achieving competitive accuracy while offering substantial computational cost savings compared to other training-free methods.
Table 2: Grounding time and token costs on Nr3D. EaSe has significant advantage, especially when compared to agent-based methods (VLM-Grounder and Transcrib3D). $\dagger$: evaluated on a subset having 250 samples.
@@
Table 3: Performances on ScanRefer. EaSe has close performance with VLM-Grounder and BUTD-DETR. $\dagger$: evaluated on a subset having 250 samples.
@@
Table 4: Performances on Nr3D. VD and VID stand for view-dependent and view-independent, respectively. We only report time and token consuming for training-free methods which involve LLM or VLM. EaSe outperforms other training-free baselines in overall performance, and has significant advantages in terms of time and token consumptions. For comparison with Transcrib3D, we use GT object labels. For comparison with NS3D, we evaluate EaSe on the same subset as NS3D. $\dagger$: VLM-Grounder is evaluated on a subset having 250 samples. * : we re-run ZSVG3D using GPT-4o.
@@ 4.3 Qualitative Results
Figure 3: Visualization of the grounding process. Anchor (the door) are marked with red circles. Objects that strongly match the conditions are highlighted in green, with brighter shades indicating higher matching scores.
@@
Figure 3 illustrates a grounding process of EaSe , demonstrating how the final grounding result is constructed through the combination of multiple conditions within the referring utterance. For example, the utterance “When facing the door, it’s the shelf above the desk on the right” can be understood as following four steps, progressing from left to right in the figure. First, objects right of the door are identified using the category feature “door” and the relation feature “right”. Next, the desk right of the door is located from this set using the category feature “desk”. Subsequently, objects satisfying the joint condition above desk and on door’s right are identified. Finally, the target “shelf” is grounded from them. More visualization results can be found in the Figure 8.
4.4 Ablation Study
Figure 4: The accuracy curves of different variants. The x-axis is the generation number of the code. The y-axis is the normalized accuracy.
We conduct ablation studies to investigate the impact of various components during the code generation and optimization processes, evaluating three different variants. We choose to analyze six relations that required multiple optimization iterations. (For relatively simple relations like “small”, the generated code passes all unit tests in the first generation, so there is no optimization.)
@@ 4.4 Ablation Study
The three variants are: Variant 1 direct prompts LLMs for multiple code, selecting the one with the highest pass rate. Variant 2 replaces the error message (see §3.3.4), which includes specific failure cases, with a general optimization instruction that omits failure cases. Variant 3 keeps all components except in-context examples. For the relations that no in-context example is used (“left”, “above”, and “corner”), variant 3 is identical to EaSe , so we only plot variant 1 and 2 on the corresponding subplots. To control for the impact of the generation results of the first iteration, we use the same results of iteration 0 across variant 2 and variant 3. In all three variants, no in-context example (§$\S$3.3.2) is used.
Figure 4 illustrates the results of the ablation study; different variants are represented by lines of different colors. The horizontal axis represents the number of iterations. The vertical axis shows the normalized accuracy on test examples associated with the relation. The effect of optimization is evident in variant 1: without optimization, LLMs fail to produce plausible spatial relation encoders for most relations, except “corner” and “between”. Variant 2 demonstrates the effect of optimization: by incorporating simple optimization, there is a noticeable performance improvement compared to variant 1. However, LLMs still struggle with relations except “above” Variant 3 highlights the effect of error messages (see § 3.3.3). By using specific failure cases in error messages, LLMs are able to generate plausible spatial relation encoders for most relations. For relations like “right”, “between” and “below” which use in-context examples, variant 3 shows a significant performance gap in initial iterations, underscoring the impact of in-context examples.
@@
In this work, we propose EaSe , a training-free 3DVG method that uses Python code to encode spatial relations and an automatic pipeline for their generation and optimization. We leverage knowledge from LLMs to create spatial relation encoders, circumventing the need for human annotation or supervised learning. EaSe eliminates the need for large-scale data and offers promising advantages in accuracy and grounding cost compared to other training-free methods.
There are some limitations in EaSe . In referring utterances, there are diverse linguistic elements beyond object categories and spatial relations. Our method, focusing primarily on these core components, may struggle with non-relation words like “second from left” in complex cases. Empirically, we observe that object bounding boxes alone are insufficient for precise spatial encoding. For simplicity, we currently ignore object shapes, orientations, and different areas within the scene such as bathroom. Integrating richer scene information into LLMs and VLMs is an important avenue for future exploration.